OpenStudy (unklerhaukus):

\[-e^{y/x} = ln (1/x) + c \] can this be solved for y?

OpenStudy (anonymous):

lets try who knows \[e^{\frac{y}{x}}=\ln(x)-c\] \[\frac{y}{x}=\ln(\ln(x)-c)\] \[y=\frac{\ln(\ln(x)-c)}{x}\]

OpenStudy (unklerhaukus):

yeah i had that but it looked kinda not right

OpenStudy (anonymous):

well i made a dumb mistake

OpenStudy (anonymous):

\[\frac{y}{x}=\ln(\ln(x)-c)\] \[y=x\ln(\ln(x)-c)\]

OpenStudy (mr.math):

satellite73 just forgot the minus sign there, which will change the expression a bit.

OpenStudy (anonymous):

hmm no i don't think i did

OpenStudy (anonymous):

i did make one stupid mistake, let me see if i made another.

OpenStudy (mr.math):

Oh I see what you did. Sorry!

OpenStudy (anonymous):

no problem, one dumb mistake is my limit!

OpenStudy (unklerhaukus):

is this likely to be a answer to a first order homogenous differential equation, i dont know if i like have the log of a log in my answer

OpenStudy (mr.math):

I knew it was a solution of a DE, what was the problem originally?

OpenStudy (unklerhaukus):

page above has working to check the question is \[xy'=y-xe^{y/x}\]

OpenStudy (mr.math):

The solution seems right to me, but I think you missed a minus sign when integrating -1/x.

OpenStudy (unklerhaukus):

so the solution has crazy logs of logs , hmm

OpenStudy (mr.math):

Yeah, crazy logs are fine too! ;D

OpenStudy (unklerhaukus):

well that is good to know thank you for your help

OpenStudy (mr.math):

No problem.